OpenStudy (tanabugg):

Write an equation in​ slope-intercept form of the line that passes through the given point and is parallel to the graph of the given equation. ​(−8​,−8​); y=−3x+5 Write an equation for the line in​ slope-intercept form.

OpenStudy (will.h):

parallel means the slope is the same thus the slope is -3 now use y - y1 = m(x - x1) y + 8 = -3(x + 8) y + 8 = -3x -24 y = -3x -32 hope that helps